# Connectedness of the Gromov boundary of fine curve graphs

### v2

@@ 3 Fine curve graph
In what follows, all curves will be essential on the surface $S$, simple, closed, and smooth, unless otherwise mentioned.
The fine curve graph $\mathcal{C}^{\dagger}(S)$ is a graph where the vertices are essential simple closed differentiable curves on $S$ and two vertices are joined by an edge if they are disjoint. The non-separating fine curve graph $\operatorname{\mathcal{NC}}^{\dagger}(S)$ is the subgraph of $\mathcal{C}^{\dagger}(S)$ consisting of vertices as non-separating curves. If they are equipped with the simplicial metrics, then these two connected graphs are quasi-isometric (with the inclusion map being a $(1,1)$-quasi-isometry) and are simultaneously Gromov hyperbolic [BHW22]. Hence by Proposition 2.3, the Gromov boundary $\partial\operatorname{\mathcal{NC}}^{\dagger}(S)$ is homeomorphic to $\partial\mathcal{C}^{\dagger}(S)$.
@@
We remark that it is possible to construct the fine curve graphs by using non-differentiable curves, i.e. $\mathcal{C}^{0}$-embedding of $S^{1}\hookrightarrow S$, which turns out to be the topological version of the fine curve graphs. However, we remark that any topological curves can be isotopic to a differentiable curve that is disjoint from it: isotopically mollify the curve and then move it away. Hence there is a $1$-Lipschitz embedding of the fine curve graphs of differentiable curves into the fine curve graphs of non-differentiable curves and this embedding is $1$-dense, i.e. the two versions of fine curve graphs are quasi-isometric and have no difference in the view point of large-scale geometry.
In the sequel, we will denote by $d^{\dagger}$ the simplicial metric on $\operatorname{\mathcal{NC}}^{\dagger}(S)$. Also, we will also denoted by $\operatorname{\mathcal{NC}}(S)$ the simplicial graph where the vertices are isotopic classes of essential non-separating simple closed curves and two vertices are connected by an edge if they admit disjoint representatives. We remark that $\operatorname{\mathcal{NC}}(S)$ is also Gromov hyperbolic and is quasi-isometric to the usual curve graph [Ham14, §3].
Definition 3.3 (Transversality).
Let $\alpha,\beta$ be two simple closed curves. We say that they are transverse if exactly one of the following holds:
  • $\alpha\cap\beta=\emptyset$.
  • For every $p\in\alpha\cap\beta$, there is an open neighbourhood $N\subset S$ of $p$ and a diffeomorphism onto the unit disk $\varphi\colon N\to D\subset\mathbb{R}^{2}$ such that $\varphi(\alpha\cap N)=(-1,1)\times\{0\}$ and $\varphi(\beta\cap N)=\{0\}\times(-1,1)$.
Note that if two curves are transverse, then by compactness, they only intersect at finitely many points.
A bigon of two transverse curves $\alpha,\beta$ in a (punctured) surface is a connected component in the complementary $\alpha\cup\beta$ that is homeomorphic to a disk and bounds exactly one subarc of $\alpha$ and one subarc of $\beta$. Two transverse curves $\alpha,\beta$ are in minimal position if they have no bigon [FM11, Proposition 1.7].
One advantage of being smooth is that the curves can be taken as geodesic under some hyperbolic metric:
Proposition 3.4 (Proposition 10, Exposé 3, [FLP91]).
Let $S$ be an orientable finite-type surface of genus at least $2$. Let $\gamma$ be an essential simple closed smooth curve. Then there exists a hyperbolic metric on $S$ such that $\gamma$ is geodesic. If in addition $\gamma$ and $\gamma^{\prime}$ are two smooth curves in minimal positions, then there exists a hyperbolic metric on $S$ such that $\gamma$ and $\gamma^{\prime}$ are simultaneously geodesic.
An important metric geometric property of the fine curve graph is that it can be approximated via the curve graph of its punctured subsurfaces:
Proposition 3.5 (Lemma 3.4 & Lemma 3.6 [BHW22]).
Suppose that $\alpha,\beta\in\operatorname{\mathcal{NC}}^{\dagger}(S)$ are transverse, and that $\alpha$ and $\beta$ are in minimal position in $S-P$, where $P\subset S$ is finite and disjoint from $\alpha\cup\beta$. Then
@@
$d_{\operatorname{\mathcal{NC}}(S-P)}([\alpha]_{S-P},[\beta]_{S-P})=d^{\dagger}(\alpha,\beta),$
where $[\alpha]_{S-P}$ and $[\beta]_{S-P}$ are respectively the isotopy classes of $\alpha$ and $\beta$ in $S-P$. Moreover, there is a geodesic $\alpha=\nu_{0},\ \nu_{1},\ \dots,\ \nu_{n}=\beta$ in $\operatorname{\mathcal{NC}}^{\dagger}(S)$ such that $\nu_{i}\cap P=\emptyset$ and $[\alpha]_{S-P}=[\nu_{0}]_{S-P},\ [\nu_{1}]_{S-P},\ \dots,\ [\nu_{n}]_{S-P}=[\beta]_{S-P}$ is also a geodesic in $\operatorname{\mathcal{NC}}(S-P)$.

### v4

@@ 3 Fine curve graph and its subgraphs
It is possible to construct the fine curve graph using non-differentiable curves, i.e., $\mathcal{C}^{0}$-embeddings of $S^{1}\hookrightarrow S$. This construction is referred to as the topological version of the fine curve graph. Notably, any topological curve can be isotoped to a differentiable curve that remains disjoint from it. This can be achieved by isotopically smoothing (or "mollifying") the curve and then moving it slightly. Consequently, there exists a $1$-Lipschitz embedding of the fine curve graph of differentiable curves into the fine curve graph of non-differentiable curves, which is also $1$-dense. Thus, the two versions of the fine curve graph are quasi-isometric, meaning they are equivalent from the perspective of large-scale geometry.
Next, we examine the relationship between the fine curve graph and the classical curve graph. For any orientable surface, including the torus, there is a natural projection from the fine curve graph to the classical curve graph.
Proposition 3.2.
Let $S$ be as above. Then the map $[\cdot]\colon(\mathcal{C}^{\dagger}(S),d^{\dagger})\to(\mathcal{C}(S),d_{S})$, which sends a simple closed curve to its isotopy class, is $1$-Lipschitz.
Proof.
Let $a,b\in\mathcal{C}^{\dagger}(S)$ be any two distinct simple closed curves on $S$. Consider a geodesic $a=x_{0},x_{1},\dots,x_{n}=b$ in $\mathcal{C}^{\dagger}(S)$ connecting $a$ to $b$. For each $i=0,\dots,n-1$, the curves $x_{i}$ and $x_{i+1}$ are disjoint. Moreover, either $x_{i}$ is isotopic to $x_{i+1}$, or $[x_{i}]\neq[x_{i+1}]$.
In the latter case, the isotopy classes $[x_{i}]$ and $[x_{i+1}]$ have disjoint representatives, specifically the curves $x_{i}$ and $x_{i+1}$. This implies that $[x_{i}]$ and $[x_{i+1}]$ are adjacent in $\mathcal{C}(S)$. Consequently, we have
$d_{S}\big{(}[x_{i}],[x_{i+1}]\big{)}\leq d^{\dagger}(x_{i},x_{i+1})$
in each case. Therefore,
$d_{S}\big{(}[a],[b]\big{)}\leq\sum_{i=0}^{n-1}d_{S}\big{(}[x_{i}],[x_{i+1}]\big{)}\leq\sum_{i=0}^{n-1}d^{\dagger}(x_{i},x_{i+1})=n=d^{\dagger}(a,b).$
This shows that the map $[\cdot]\colon\mathcal{C}^{\dagger}(S)\to\mathcal{C}(S)$ is $1$-Lipschitz. ∎
Suppose now that $S$ is equipped with a hyperbolic metric. Then each simple closed curve on $S$ is isotopic to a unique geodesic representative [farb2011primer, Proposition 1.3]. Thus, the map $\gamma\colon\mathcal{C}(S)\hookrightarrow\mathcal{C}^{\dagger}(S)$, defined by sending an isotopy class $a\in\mathcal{C}(S)$ to its geodesic representative $\gamma(a)\in\mathcal{C}^{\dagger}(S)$, is well-defined. A similar argument to that in Proposition 3.2 shows the following result.
Proposition 3.3.
Let $S$ be a hyperbolic surface. Then the map $\gamma\colon\mathcal{C}(S)\hookrightarrow\mathcal{C}^{\dagger}(S)$, which assigns to each isotopy class $a\in\mathcal{C}(S)$ its unique geodesic representative $\gamma(a)\in\mathcal{C}^{\dagger}(S)$, is an isometric embedding.
Proof.
Let $\alpha,\beta\in\mathcal{C}(S)$, and let $\alpha=y_{0},\dots,y_{m}=\beta$ be a geodesic in $\mathcal{C}(S)$ connecting $\alpha$ to $\beta$. For each $i=0,\dots,m-1$, the isotopy classes $y_{i}$ and $y_{i+1}$ admit disjoint representatives. Since geodesics are in minimal position [farb2011primer, Corollary 1.9], it follows that the geodesic representatives $\gamma(y_{i})$ and $\gamma(y_{i+1})$ must also be disjoint. Thus, we have $d^{\dagger}\big{(}\gamma(y_{i}),\gamma(y_{i+1})\big{)}=1$ for each $i$.
As a result,
$d^{\dagger}\big{(}\gamma(\alpha),\gamma(\beta)\big{)}\leq\sum_{i=0}^{m-1}d^{\dagger}\big{(}\gamma(y_{i}),\gamma(y_{i+1})\big{)}=m=d_{S}(\alpha,\beta),$
showing that $\gamma$ is $1$-Lipschitz.
Now, note that the composition $[\cdot]\circ\gamma=\mathrm{Id}\colon\mathcal{C}(S)\to\mathcal{C}(S)$ is the identity map. By Proposition 3.2, we have
$d_{S}(\alpha,\beta)=d_{S}\big{(}[\gamma(\alpha)],[\gamma(\beta)]\big{)}\leq d^{\dagger}\big{(}\gamma(\alpha),\gamma(\beta)\big{)}.$
Therefore, $d_{S}(\alpha,\beta)=d^{\dagger}\big{(}\gamma(\alpha),\gamma(\beta)\big{)}$, which shows that $\gamma$ is an isometry.
Finally, it is clear that $\gamma$ is also an embedding, completing the proof. ∎
3.B. Distance estimation.
Similar to the curve graph, we can estimate distances between certain curves by examining the cardinality of their intersections. To proceed, we introduce the following concept:
Definition 3.4 (Transversality).
Let $\alpha,\beta$ be two simple closed curves on $S$. We say that they are transverse if exactly one of the following holds:
  • $\alpha\cap\beta=\emptyset$;
  • For each point $p\in\alpha\cap\beta$, there exists an open neighbourhood $N\subset S$ of $p$ and a diffeomorphism $\varphi\colon N\to D\subset\mathbb{R}^{2}$ onto the unit disc $D$, such that $\varphi(\alpha\cap N)=(-1,1)\times\{0\}$ and $\varphi(\beta\cap N)=\{0\}\times(-1,1)$.
Note that if two curves are transverse, then by the compactness of the surface, they intersect at only finitely many points.
A bigon between two transverse curves $\alpha$ and $\beta$ on a surface is a connected component of the complement $S\setminus(\alpha\cup\beta)$ that is homeomorphic to a disc and contains exactly one subarc of $\alpha$ and one subarc of $\beta$. Two transverse curves $\alpha$ and $\beta$ are said to be in minimal position if they contain no bigon [farb2011primer, Proposition 1.7].
One advantage of smoothness is that curves can be taken as geodesics under an appropriate hyperbolic metric:
Proposition 3.5 (Proposition 10, Exposé 3, [fathi1991travaux]).
Let $S$ be a connected orientable surface of genus $g\geq 2$. Let $\gamma$ be a curve on $S$. Then there exists a hyperbolic metric on $S$ such that $\gamma$ is geodesic. Moreover, if $\gamma$ and $\gamma^{\prime}$ are two smooth curves in minimal position, then there exists a hyperbolic metric on $S$ such that both $\gamma$ and $\gamma^{\prime}$ are geodesic.
In particular, by combining Proposition 3.5 and Proposition 3.3, we obtain the following result:
Corollary 3.6.
For any two vertices $\alpha,\beta\in\mathcal{C}^{\dagger}(S)$ that are in minimal position on $S$, we have $d_{S}([\alpha],[\beta])=d^{\dagger}(\alpha,\beta)$, where $d_{S}$ denotes the simplicial distance in the curve graph $\mathcal{C}(S)$.
Now, let $P\subset S$ be a finite collection of points. Adapting terminology from [bowden2022quasi], recall that a curve $\gamma\in\mathcal{C}^{\dagger}(S-P)$ is called surviving if it remains essential in $S$. We denote by $\mathcal{C}^{s}(S-P)$ the surviving curve graph of the punctured surface $S-P$. An important property in the metric geometry of the fine curve graph is that it can be approximated by the surviving curve graphs of these punctured subsurfaces:
Proposition 3.7 ([bowden2022quasi]).
Suppose that $\alpha,\beta\in\mathcal{C}^{\dagger}(S)$ are transverse, and that $\alpha$ and $\beta$ are in minimal position in $S-P$, where $P\subset S$ is finite and disjoint from $\alpha\cup\beta$. Then
$d^{\dagger}(\alpha,\beta)=d_{\mathcal{C}^{s}(S-P)}\left([\alpha]_{S-P},[\beta]_{S-P}\right)$
where $[\alpha]_{S-P}$ and $[\beta]_{S-P}$ are respectively the isotopy classes of $\alpha$ and $\beta$ in $S-P$. Moreover, if $\alpha$ and $\beta$ are two non-separating curves, then
$d^{\dagger}(\alpha,\beta)=d_{\mathcal{NC}(S-P)}([\alpha]_{S-P},[\beta]_{S-P}),$
and that the inclusion map $\mathcal{NC}^{\dagger}(S)\hookrightarrow\mathcal{C}^{\dagger}(S)$ is a $(1,2)$-quasi-isometry.
Similar to [masur1999geometry, Lemma 2.1], we can deduce the following distance estimation in the fine curve graph:
Proposition 3.8.
Let $S$ be a surface with genus $g(S)\geq 2$. For any two transverse curves $\alpha,\beta\in\mathcal{C}^{\dagger}(S)$, we have $d^{\dagger}(\alpha,\beta)\leq 2|\alpha\cap\beta|+1$.
Proof.
If $|\alpha\cap\beta|=0$, then the formula is trivially true by the definition of the simplicial distance in $\mathcal{C}^{\dagger}(S)$.
If $|\alpha\cap\beta|=1$, then there is no bigon between $\alpha$ and $\beta$, meaning that these two curves are in minimal position. Since $g(S)\geq 2$, the subsurface $S-(\alpha\cap\beta)$ has positive genus, which implies that we can find a non-separating (and thus surviving) curve disjoint from both $\alpha$ and $\beta$, so that $d^{\dagger}(\alpha,\beta)=2$.
Now suppose $|\alpha\cap\beta|\geq 2$. Let $P\subset S$ be a finite set of points such that $\alpha$ and $\beta$ are in minimal position on $S-P$. We can perform the same surgery on $S-P$ for $\beta$ as in [masur1999geometry, Lemma 2.1]. Note that we will have two possibilities, $\beta_{1}$ and $\beta_{2}$, obtained by replacing a subarc of $\beta$ with a segment isotopic to a subarc of $\alpha$. Both $\beta_{1}$ and $\beta_{2}$ remain essential on $S-P$ by [masur1999geometry, Lemma 2.1]. It suffices to show that the resulting curve is a surviving curve. If both $\beta_{1}$ and $\beta_{2}$ are non-surviving, then so will $\beta$, which contradicts the assumption that $\beta\in\mathcal{C}^{\dagger}(S)$ and survives on $S-P$.
By induction, we can then find a path of length $12|\alpha\cap\beta|+$ 1 connecting $[\alpha]_{S-P}$ to $[\beta]_{S-P}$ in the surviving curve graph $\mathcal{C}^{s}(S-P)$, which leads to the desired distance estimate in $\mathcal{C}^{\dagger}(S)$ after applying Proposition 3.7. ∎
Recall that a regular neighbourhood $\mathcal{N}$ of a simple closed curve $\gamma$ in $S$ is a locally flat, compact submanifold of $S$ of codimension 00, which is a topological neighbourhood of $\gamma$ such that the inclusion $\gamma\subset\mathcal{N}$ is a simple homotopy equivalence, and $\gamma$ is a strong deformation retract of $\mathcal{N}$. We note that, by definition, this regular neighbourhood can be chosen arbitrarily small. Furthermore, if the surface is endowed with a metric, then any annular neighbourhood with sufficiently small width will be a regular neighbourhood.
Lemma 3.9.
Let $\alpha\in\mathcal{C}^{\dagger}(S)$ and let $F\subset\mathcal{C}^{\dagger}(S)$ be a finite collection of pairwise transverse curves. Then there exists a curve $\alpha^{\prime}\in\mathcal{C}^{\dagger}(S)$, isotopic to $\alpha$, that is disjoint from $\alpha$ and transverse to all curves in $F$.
Proof.
Define
$C\coloneqq\{x\in S:x\in\gamma_{i}\cap\gamma_{j}\text{ for some }\gamma_{i},\gamma_{j}\in F\text{ and }i\neq j\}.$
Let $N(\alpha)$ be a sufficiently small regular neighbourhood of $\alpha$. Consider a small enough regular neighbourhood $N(C)$ of $C$ and a sufficiently small neighbourhood $N_{i}$ of each $\gamma_{i}$, such that $N_{i}\cap N_{j}\subset N(C)$ whenever $i\neq j$. First, isotope $\alpha$ to a curve $\alpha_{0}$ away from $N(C)\cup N(\alpha)$. Note that any isotopy supported within $N_{i}-N(C)$ will not affect the intersection behaviour between $\alpha_{0}$ and $\gamma_{j}$ for all $j\neq i$. Hence, we can isotope $\alpha_{0}$ within $N_{i}-\big{(}N(C)\cup N(\alpha)\big{)}$ for $i=1,\dots,|F|$. In this way, we can find a curve $\alpha^{\prime}$ isotopic to $\alpha$, disjoint from $\alpha$, and transverse to all $\gamma_{i}\in F$. ∎
@@
We note that the perturbation in Lemma 3.9 can be performed within an arbitrarily small regular neighbourhood. All isotopy perturbations in the sequel will follow the same basic method, although the specific situation might vary slightly.
3.C. Non-separating subgraph.
The non-separating fine curve graph $\operatorname{\mathcal{NC}}^{\dagger}(S)$ is the subgraph of $\mathcal{C}^{\dagger}(S)$ consisting of vertices corresponding to non-separating curves. When equipped with the simplicial metric, these two connected graphs are quasi-isometric, with the inclusion map being a $(1,1)$-quasi-isometry, and both graphs are simultaneously Gromov hyperbolic. Hence, by Proposition 2.3, the Gromov boundary $\partial\operatorname{\mathcal{NC}}^{\dagger}(S)$ is homeomorphic to $\partial\mathcal{C}^{\dagger}(S)$.
To make full use of the known results about the classical non-separating curve graph, let us fix an arbitrary base point $o\in\mathcal{NC}^{\dagger}(S)$ and denote by $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ the subgraph of $\mathcal{NC}^{\dagger}(S)$ consisting of vertices that are all transverse to $o$. This subgraph inherits the subgraph distance $d_{o}^{\dagger}$, but not the induced distance $d^{\dagger}$ from $\operatorname{\mathcal{NC}}^{\dagger}(S)$.
Proposition 3.11.
The inclusion map $i:\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\hookrightarrow\operatorname{\mathcal{NC}}^{\dagger}(S)$ is a $(1,2)$-quasi-isometry, hence $\partial\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\simeq\partial\operatorname{\mathcal{NC}}^{\dagger}(S)$.
Proof.
We first show that $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ is $1$-dense in $\operatorname{\mathcal{NC}}^{\dagger}(S)$. Consider any $\alpha\in\operatorname{\mathcal{NC}}^{\dagger}(S)$. If $\alpha$ is not transverse to $o$, then, by Lemma 3.9, there exists another curve $\alpha^{\prime}\in\mathcal{NC}^{\dagger}(S)$ that is disjoint from $\alpha$ and is transverse to $o$. Consequently, we have
@@
$d^{\dagger}\left(\alpha,\alpha^{\prime}\right)=d^{\dagger}\left(\alpha,\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\right)=1.$
This shows that the inclusion map is essentially surjective.
If $\alpha,\beta\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ are transverse, then by taking a finite subset $P\subset S$ such that $o,\alpha,\beta$ are pairwise in minimal position on $S-P$, Proposition 3.7 indicates that
$d^{\dagger}_{o}(\alpha,\beta)=d^{\dagger}(\alpha,\beta)=d_{\mathcal{N}\mathcal{C}(S-P)}\left([\alpha]_{S-P},[\beta]_{S-P}\right).$
Now for an arbitrary pair $\alpha,\beta\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, we can perturb them to a transverse pair, which will cause an additive error at most $2$. This implies that the inclusion map is a $(1,2)$-quasi-isometry. As a result, their Gromov boundaries are homeomorphic (see Proposition 2.3). ∎
Finally, we need to point out the following fact about the comparison between two distances $d^{\dagger}$ and $d^{\dagger}_{o}$:
Lemma 3.12.
For any $\alpha\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, we always have $d_{o}^{\dagger}(o,\alpha)=d^{\dagger}(o,\alpha)$.
Proof.
Indeed, for any $\alpha\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, it is clear that the subgraph distance is always bounded from above by the induced distance, i.e. $d^{\dagger}(\alpha,o)\leq d_{o}^{\dagger}(\alpha,o)$. Conversely, because $\alpha$ and $o$ are transverse, by Proposition 3.7, there exists a finite subset $P\subset S$ disjoint from $o\cup\alpha$ such that $o$ and $\alpha$ are in minimal position in $S-P$. Moreover, we have
$d_{\operatorname{\mathcal{NC}}(S-P)}([o]_{S-P},[\alpha]_{S-P})=d^{\dagger}(o,\alpha).$
Let $[o]_{S-P}=[\nu_{0}]_{S-P},\ [\nu_{1}]_{S-P},\ \dots,\ [\nu_{n}]_{S-P}=[\alpha]_{S-P}$ be a geodesic in $\operatorname{\mathcal{NC}}(S-P)$. By Proposition 3.5, we can choose a hyperbolic metric on $S-P$ so that $o$ and $\alpha$ are simultaneously geodesic. Let $\nu_{i}$ be the geodesic representative of $[\nu_{i}]_{S_{P}}$. Now we can find a path
$o=\nu_{0},\nu_{1},\cdots,\nu_{n}=\alpha,$
connecting $o$ to $\alpha$ in $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$. This implies that $d_{o}^{\dagger}(o,\alpha)\leq n=d^{\dagger}(o,\alpha)$. This proves the desired result. ∎